# Lifetime-Aware Design of Item-Level Intelligence

### v1

@@ Abstract.
We present FlexiFlow, a lifetime-aware design framework for item-level intelligence (ILI) where computation is integrated directly into disposable products like food packaging and medical patches. Our framework leverages natively flexible electronics which offer significantly lower costs than silicon but are limited to kHz speeds and several thousands of gates. Our insight is that unlike traditional computing with more uniform deployment patterns, ILI applications exhibit $1000\times$ variation in operational lifetime, fundamentally changing optimal architectural design decisions when considering trillion-item deployment scales. To enable holistic design and optimization, we model the trade-offs between embodied carbon footprint and operational carbon footprint based on application-specific lifetimes. The framework includes: (1) FlexiBench, a workload suite targeting sustainability applications from spoilage detection to health monitoring; (2) FlexiBits, area-optimized RISC-V cores with 1/4/8-bit datapaths achieving $2.65\times$ to $3.50\times$ better energy efficiency per workload execution; and (3) a carbon-aware model that selects optimal architectures based on deployment characteristics. We show that lifetime-aware microarchitectural design can reduce carbon footprint by $1.62\times$, while algorithmic decisions can reduce carbon footprint by $14.5\times$. We validate our approach through the first tape-out using a PDK for flexible electronics with fully open-source tools, achieving 30.9 kHz operation. FlexiFlow enables exploration of computing at the Extreme Edge where conventional design methodologies must be reevaluated to account for new constraints and considerations.
@@ 1. Introduction
FlexiBench provides the first benchmark suite of workloads specifically designed for ILI applications. Unlike traditional benchmarks in embedded computing that focus on performance and energy efficiency, FlexiBench comprises 11 workloads targeting real-world sustainability challenges aligned with UN SDGs. These workloads span diverse computational patterns (simple comparison checks to tiny machine learning inference), memory requirements (0.3 KB to 240 KB), and expected usage lifetime (days to years) and frequency (real-time to daily).
@@
2. Background, Motivation, & Related Work
@@ 2.3. Related Work
Design Methodologies. While some frameworks have been proposed for flexible electronics (Raisiardali et al., 2025; Ozer et al., 2024c; Bleier et al., 2022; Ozer et al., 2023), the work is typically focused on traditional PPA metrics. As we argue in Section 4.4, these metrics often do not directly affect on workload success and crucially do not take into account the sustainability challenges of ILI.
Our work addresses gaps in all three infrastructure components: FlexiBench provides a set of sustainability-focused workloads, FlexiBits offers multiple processor implementations using a standard 32-bit ISA, and FlexiFlow connects them through PPA-driven carbon-optimal selection. To further advance progress towards ILI, we open-source our work at https://bit.ly/FlexiFlow.
@@ 3. FlexiBench
Medium-term Deployments (Months)
@@
Long-term Deployments (Years)
@@ 4.3. Flexible Bitwidth Architectures
FlexiBits RISPs FlexiCores PlasticARM
@@
Table 4. Comparison of recent flexible processors. FlexiBits uniquely offers multiple area-efficient implementations.
@@ 4.4. Performance, Power, and Area Comparison
Tape-Out Performance. To further validate our designs, we tape-out a full SERV-based SoC using the open-source OpenROAD (Ajayi et al., 2019) on a FlexIC. Implementation achieved a frequency of 30.9 kHz, and all fabricated dies reliably operated at 33.0 kHz. Further discussion of our tape-out in the broader context of FlexiFlow are in Section 6.5 and Appendix B.4.
@@ 5.2. Framework Inputs
Foundry Data. Lastly, FlexiFlow incorporates foundry data, including the Process Development Kit (PDK) for physical implementation and LCA that quantifies the embodied carbon footprint associated with fabrication. For our studies, we use Pragmatic Semiconductor’s PDK and standard cell library for developing FlexICs along with their embodied carbon metrics based on a proprietary analytical model. While FlexiFlow is agnostic to a specific technology in its structure, we reiterate that ILI requires form factors, cost profiles, and carbon footprints that are incompatible with traditional silicon-based technologies. Thus, flexible electronics offer a more viable pathway and is the focus of our framework.
@@ 5.5. Lifetime-Aware Design Model
Figure 5. Carbon-optimal system selection for ILI applications generated using FlexiFlow depending on lifetime and program execution frequency. Red stars correspond to the example use cases from Table 2. Tree tracking workload omitted due to extremely high task compute time.
@@ 6.2. Lifetime-Aware Architecture Selection
What Drives These Trade-offs? The inflection points in Figure 5 depend on workload characteristics. VM-intensive and high-work workloads like Gesture Recognition see earlier transitions to efficient processors. Their high SRAM power consumption and long individual task execution times amplify the benefits of faster processing. Conversely, simple workloads like Malodor Classification favor SERV, as their minimal work generates negligible operational differences.
@@ 6.4. At Scale Computing for Sustainability
It should be noted that many computing for sustainability applications, including many in FlexiBench, cannot be quantified by an equivalent carbon footprint. Furthermore, the benefits of food spoilage detection go beyond equivalent carbon savings: Saving food minimizes environmental impacts across critical planetary boundaries (Richardson et al., 2023) (e.g., land and water use), enhances food security, and reduces waste management burdens. We present this scaling study only to illustrate a quantifiable benefit of using ILI for the UN SDGs.
@@ 6.5. Open-Source for ILI with Tape-Out Validation
(a) Die shot of FlexIC.
(b) Die assembled on FlexPCB.
Figure 7. Open-source tape out from FlexiFlow.
@@ 7. Conclusion
Item-level intelligence (ILI) presents new opportunities to deeply embed computing beyond where it has traditionally existed. However, ILI has yet to materialize in part due to the absence of open-source research and development infrastructure necessary to enable systematic exploration and progress. Our end-to-end framework with open-source benchmarks and microprocessors aims to take one step towards enabling ILI with flexible electronics, while balancing computing for sustainability opportunities and sustainable computing considerations. FlexiFlow provides insights and design strategies for this emerging domain to inform designers of the architectural and algorithmic trade-offs to be considered at the item level. Our framework lays key foundations for future work to continue advancing tools and technologies that enable the widespread realization of ILI at scale.
@@ B.1. Performance, Power, and Area Detailed Results
Exact area and power numbers for the FlexiBits processors are in Table 7.
@@
Memory. In Table 8, we characterize the power requirements of system SRAM and LPROM based on the profiled memory requirements of each workload in Section 3 (Table  3). In this profile, we find SRAM dominates the memory power footprint, while LPROM contributes negligibly to total power consumption. Consequently, workloads with higher dynamic memory requirements exhibit greater overall memory power. We further note that any contribution to total power by SRAM will only amplify the energy benefit provided by a faster core. Therefore, workloads with higher memory power consumption, such as package tracking and gesture recognition (Figure 10), will benefit greatly from higher performance cores. As it stands today, the memory requirements of some benchmarks exceed what is currently supported by today’s FlexIC technology (e.g., gesture recognition). We leave design and optimization of the memory subsystem to future work. Embedded DRAM leveraging IGZO could enable random access memory with improved energy-efficiency, area footprint and carbon footprint (Kong et al., 2025; Belmonte et al., 2020; Grey-Stewart et al., 2025).
@@
Figure 12. Simulated effect of instruction mixes at the two extremes in the FlexiBits microarchitecture. The workload on the left-hand plot comprises only one-stage instructions. (right) The workload on the right-hand side contains only two-stage instructions. Marginal differences in the inflection points between optimal microarchitectures can be observed across the two simulated instruction mixes.
@@ B.3.1. Instruction Mix
Notably, our cores do differentiate between one-stage and two-stage instructions. A workload that contains all two-stage instructions will take about $2\times$ longer than a workload that contains only one-stage instructions (assuming a constant number of instructions). In this extreme, a “two-stage only” workload would place inflection points at marginally earlier lifetimes and task frequencies. A visualization of this can be seen in Appendix Figure 12. We place an example point close to the border between QERV and HERV, showing how instruction mix could change optimal design in extreme edge cases.
@@ B.3.2. Energy Source
Figure 13 visualizes how this parameter affects carbon-optimal design decisions for the Air Pollution Monitoring workload. In that particular workload, the optimal microarchitecture design changes with energy source due to these differences. Therefore, the expected energy source is an important user-specification as described in FlexiFlow’s inputs (Section 5.2).

### v2

# Lifetime-Aware Design for Item-Level Intelligence at the Extreme Edge

@@ Abstract.
We present FlexiFlow, a lifetime-aware design framework for item-level intelligence (ILI) where computation is integrated directly into disposable products like food packaging and medical patches. Our framework leverages natively flexible electronics which offer significantly lower costs than silicon but are limited to kHz speeds and several thousands of gates. Our insight is that unlike traditional computing with more uniform deployment patterns, ILI applications exhibit $1000\times$ variation in operational lifetime, fundamentally changing optimal architectural design decisions when considering trillion-item deployment scales. To enable holistic design and optimization, we model the trade-offs between embodied carbon footprint and operational carbon footprint based on application-specific lifetimes. The framework includes: (1) FlexiBench, a workload suite targeting sustainability applications from spoilage detection to health monitoring; (2) FlexiBits, area-optimized RISC-V cores with 1/4/8-bit datapaths achieving $2.65\times$ to $3.50\times$ better energy efficiency per workload execution; and (3) a carbon-aware model that selects optimal architectures based on deployment characteristics. We show that lifetime-aware microarchitectural design can reduce carbon footprint by $1.62\times$, while algorithmic decisions can reduce carbon footprint by $14.5\times$. We validate our approach through the first tape-out using a PDK for flexible electronics with fully open-source tools, achieving 30.9 kHz operation. FlexiFlow enables exploration of computing at the Extreme Edge where conventional design methodologies must be reevaluated to account for new constraints and considerations. FlexiFlow is available at https://github.com/harvard-edge/FlexiFlow.
Extreme Edge; TinyML; Sustainability; Carbon Footprint; RISC-V; Printed and Flexible Electronics; FlexIC; System Analysis and Design
journalyear: 2026copyright: ccconference: Proceedings of the 31st ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2; March 22–26, 2026; Pittsburgh, PA, USAbooktitle: Proceedings of the 31st ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2 (ASPLOS ’26), March 22–26, 2026, Pittsburgh, PA, USAdoi: 10.1145/3779212.3790182isbn: 979-8-4007-2359-9/2026/03ccs: Hardware Analysis and design of emerging devices and systemsccs: Computer systems organization Embedded systemsccs: Hardware Impact on the environmentccs: Human-centered computing Ubiquitous and mobile computingccs: Social and professional topics Sustainability
@@ 1. Introduction
FlexiBench provides the first benchmark suite of workloads specifically designed for ILI applications. Unlike traditional benchmarks in embedded computing that focus on performance and energy efficiency, FlexiBench comprises 11 workloads targeting real-world sustainability challenges aligned with the UN SDGs. These workloads span diverse computational patterns (simple comparison checks to tiny machine learning inference), memory requirements (0.3 KB to 240 KB), and expected usage lifetime (days to years) and frequency (real-time to daily).
@@
2. Background, Motivation & Related Work
@@ 2.3. Related Work
Design Methodologies. While some frameworks have been proposed for flexible electronics (Raisiardali et al., 2025; Ozer et al., 2024c; Bleier et al., 2022; Ozer et al., 2023), the work is typically focused on traditional PPA metrics. As we argue in Section 4.4, these metrics often do not directly affect workload success and crucially do not take into account the sustainability challenges of ILI.
Our work addresses gaps in all three infrastructure components: FlexiBench provides a set of sustainability-focused workloads, FlexiBits offers multiple processor implementations using a standard 32-bit ISA, and FlexiFlow connects them through PPA-driven carbon-optimal selection. To further advance progress towards ILI, we open-source our work at https://github.com/harvard-edge/FlexiFlow.
@@ 3. FlexiBench
Medium-lived Deployments (Months)
@@
Long-lived Deployments (Years)
@@ 4.3. Flexible Bitwidth Architectures
FlexiBits RISSPs FlexiCores PlasticARM
@@
Table 4. Comparison of recent flexible processors. FlexiBits uniquely offers multiple area-efficient implementations. SoC denotes designs with integrated CPU, on-chip memory, and peripheral interfaces.
In this work, SERV, QERV, and HERV are three representative instantiations of our design approach (Figure 3). They demonstrate that, even under tight area constraints, treating datapath width as a first-class, tunable parameter yields a rich design space with significant area-energy trade-offs (Section 4.4). Automation to exhaustively enumerate and generate all datapath widths with our template is left to future work.
@@ 4.4. Performance, Power, and Area Comparison
Tape-Out Performance. To further validate our designs, we tape-out a full SERV-based SoC using the open-source OpenROAD (Ajayi et al., 2019) on a FlexIC. Implementation achieved a frequency of 30.9 kHz, and fabricated dies reliably operated at 33.0 kHz. Further discussion of our tape-out in the broader context of FlexiFlow are in Section 6.5 and Appendix B.4.
@@ 5.2. Framework Inputs
Foundry Data. Lastly, FlexiFlow incorporates foundry data, including the Process Development Kit (PDK) for physical implementation and LCA that quantifies the embodied carbon footprint associated with fabrication. For our studies, we use Pragmatic Semiconductor’s PDK and standard cell library for developing FlexICs along with their embodied carbon metrics based on a proprietary analytical model. While FlexiFlow is agnostic to a specific technology in its structure, we reiterate that ILI requires form factors, cost profiles, and carbon footprints that are incompatible with traditional silicon-based technologies. Thus, flexible electronics offer a more viable pathway and are the focus of our framework.
@@ 5.4. Sustainability Metrics
Figure 5. Carbon-optimal system selection for ILI applications generated using FlexiFlow depending on lifetime and program execution frequency. Red stars correspond to the example use cases from Table 2 and are accompanied with per-device total carbon footprints. Tree tracking workload omitted due to extremely high task compute time.
@@ 6.2. Lifetime-Aware Architecture Selection
What Drives These Trade-offs? The inflection points in Figure 5 depend on workload characteristics. VM-intensive and high-work workloads like Gesture Recognition see earlier transitions to efficient processors. Their high SRAM power consumption and long individual task execution times amplify the benefits of faster processing. Conversely, simple workloads like Malodor Classification favor SERV as their minimal work generates negligible operational differences.
@@ 6.4. At Scale Computing for Sustainability
(a) Die shot of FlexIC.
(b) Die assembled on FlexPCB.
Figure 7. Open-source tape out from FlexiFlow.
@@
It should be noted that many computing for sustainability applications, including many in FlexiBench, cannot be quantified by an equivalent carbon footprint. Furthermore, the benefits of food spoilage detection go beyond equivalent carbon savings: Saving food minimizes environmental impacts across critical planetary boundaries (Richardson et al., 2023) (e.g., land and water use), enhances food security, and reduces waste management burdens.
We present this scaling study only to illustrate a quantifiable benefit of using ILI for the UN SDGs. Additional considerations, such as material waste, should be incorporated as more comprehensive public LCA data becomes available. While our analysis suggests that ILI can yield a net reduction in overall waste, mass deployment of FlexICs will introduce new waste streams and therefore warrants careful consideration. Promising research on this topic of e-waste is growing (Jain et al., 2025) with prior work identifying viable end-of-life pathways for FlexIC-enabled plastic packaging (Ahamed et al., 2024). Moreover, FlexICs can promote better product renewal and recycling by enabling digital product passports (European Union, 2024). Nevertheless, scaling to trillions of devices requires ongoing scrutiny of these trade-offs. To support such assessments, FlexiFlow can be naturally extended to incorporate material-waste metrics alongside carbon accounting data.
@@ 7. Conclusion
Item-level intelligence (ILI) presents new opportunities to deeply embed computing beyond where it has traditionally existed. However, ILI has yet to materialize in part due to the absence of open-source research and development infrastructure necessary to enable systematic exploration and progress. Our end-to-end framework with open-source benchmarks and microprocessors aims to take one step towards enabling ILI with flexible electronics, while balancing computing for sustainability opportunities and sustainable computing considerations. FlexiFlow provides insights and design strategies for this emerging domain to inform designers of the architectural and algorithmic trade-offs that should be considered at the item-level. Our framework lays key foundations for future work to continue advancing tools and technologies that enable the widespread realization of ILI at scale.
Acknowledgements.
The authors would like to thank the anonymous reviewers for their feedback on improving the manuscript’s quality. This work was supported by NSF Grant CCF-2324862. The authors were additionally supported by the NSF Graduate Research Fellowship Program (GRFP).
@@ B.1. Performance, Power, and Area Detailed Results
[…]
Memory. In Table 8, we characterize the power requirements of system SRAM and LPROM based on the profiled memory requirements of each workload in Section 3 (Table  3). In this profile, we find SRAM dominates the memory power footprint, while LPROM contributes negligibly to total power consumption. Consequently, workloads with higher dynamic memory requirements exhibit greater overall memory power. We further note that any contribution to total power by SRAM will only amplify the energy benefit provided by a faster core. Therefore, workloads with higher memory power consumption, such as package tracking and gesture recognition (Figure 10), will benefit greatly from higher performance cores. As it stands today, the memory requirements of some benchmarks exceed what is currently supported by today’s FlexIC technology (e.g., gesture recognition). We leave design and optimization of the memory subsystem to future work. Embedded DRAM leveraging IGZO could enable random access memory with improved energy-efficiency, area footprint, and carbon footprint (Kong et al., 2025; Belmonte et al., 2020; Grey-Stewart et al., 2025).
@@ B.3.1. Instruction Mix
Notably, our cores do differentiate between one-stage and two-stage instructions. A workload that contains all two-stage instructions will take about $2\times$ longer than a workload that contains only one-stage instructions (assuming a constant number of instructions). In this extreme, a “two-stage only” workload would place inflection points at marginally earlier lifetimes and task frequencies. A visualization of this can be seen in Appendix Figure 12. We place an example point close to the border between QERV and HERV, showing how instruction mix could change optimal design in unique edge cases.
@@ B.3.2. Energy Source
Figure 12. Simulated effect of instruction mixes at the two extremes in the FlexiBits microarchitecture. The workload on the left-hand plot comprises only one-stage instructions. The workload on the right-hand side contains only two-stage instructions. Marginal differences in the inflection points between optimal microarchitectures can be observed across the two simulated instruction mixes.
@@
Figure 13 visualizes how this parameter affects carbon-optimal design decisions for the Air Pollution Monitoring workload. For this particular workload, the optimal microarchitecture design changes with energy source due to these differences. Therefore, the expected energy source is an important user-specification as described in FlexiFlow’s inputs (Section 5.2).